OpenStudy (anonymous):

Determine the quadratic function f(x) = ax^2+bx+c, whose vertex is (-2, -11) and passes through the point (-5, 7). Please show all of your work. @Calculus1

OpenStudy (wasiqss):

listen make factors into (x+2) and (x+5), because at these points of x equation would be valid and posses a certain value of y, so then multiply these factors to get the answer

OpenStudy (wasiqss):

and after making equation if u put these value of x that is -2,-5, then it should the y as(-11,7)

OpenStudy (wasiqss):

medals needed :P

OpenStudy (anonymous):

hoose the least common multiple of 12 and 24. A. 6 B. 12 C. 24 D. 36

OpenStudy (anonymous):

(x+2)^2 -(1/2)(y+11)=0 eq of parabola

OpenStudy (anonymous):

how did you get this answer Mark O. ??

OpenStudy (anonymous):

can you please show the work..thanks

OpenStudy (anonymous):

vertex is (-2, -11) use, 4a(y--11)=(x--2)^2 4a(y+11)=(x+2)^2,passes through the point (-5, 7). 4a(7+11)=(-5+2)^2 4a(18)=(-3)^2 72a=9 a=9/72=1/8, therefore 4(1/8)(y+11)=(x+2)^2 (1/2)(y+11)=(x+2)^2 (x+2)^2 -(1/2)(y+11)=0 eq of parabola

OpenStudy (anonymous):

ask me if you have question

OpenStudy (anonymous):

thank you mark o. I really appreciate it. :)